# Some Results on Finitely Splitting Subtrees of Aronszajn Trees

### v1

@@ 7. Not Adding Small-Splitting Subtrees
Let $1<n\leq s$ and suppose that $S$ is an $\omega_{1}$-tree which does not contain an uncountable downwards closed $(<\!n)$-splitting subtree. Consider a condition $(F,\Gamma)\in\mathbb{P}$ with some top level $\beta$ and assume that $(F,\Gamma)$ forces that $\dot{W}$ is a downwards closed $(<\!n)$-splitting subtree of $S$. We find an extension of $(F,\Gamma)$ which forces that $\dot{W}$ is countable.
@@ 8. Preserving Suslin Trees
Define $B$ as the set of all $b\in S$ satisfying:
  • $c\upharpoonright\beta<_{S}b$;
  • letting $\delta=\mathrm{ht}_{S}(b)$, there exists some $\vec{y}=(y_{0},\ldots,y_{l-1})\in T_{\vec{x}}\cap T_{\delta}^{l}$ satisfying the next statement;
  • whenever $(G,\Sigma)\leq(F,\Gamma)$ has some height $\gamma\leq\delta$ and for all $i<l$, $G(y_{i}\upharpoonright\gamma)=1$, then it is not the case that $(G,\Sigma)\Vdash b\in\dot{E}$.
@@
Note that these properties imply that $B$ is uncountable and downwards closed in $S_{c\upharpoonright\beta}$.
For any $e\in B$, let $\delta_{e}=\mathrm{ht}_{S}(e)$ and define $\mathcal{Y}_{e}$ to be the set of all $\vec{y}=(y_{0},\ldots,y_{l-1})\in T_{\vec{x}}\cap T_{\delta_{e}}^{l}$ witnessing that $e\in B$, which means that whenever $(G,\Sigma)\leq(F,\Gamma)$ has some height $\gamma\leq\delta_{e}$ and for all $i<l$, $G(y_{i}\upharpoonright\gamma)=1$, then it is not the case that $(G,\Sigma)\Vdash e\in\dot{E}$. Note that by (3) above, for all $e$ and $f$ in $B$ such that $e<_{S}f$, $\mathcal{Y}_{f}\upharpoonright\delta_{e}\subseteq\mathcal{Y}_{e}$.
@@
By our assumption on $S$, $T$ is still Aronszajn in $V[G_{S}]$. So we can apply Lemma 5.6 in $V[G_{S}]$ to find a countable limit ordinal $\delta>\beta$ such that the function $\mathcal{U}$ with domain $[\delta,\omega_{1})$ defined by $\mathcal{U}(\gamma)=\{U\cap T_{\gamma}^{l}\}$ is a generalized promise which is suitable for $F$. Now let $\dot{W}$, $\dot{U}$, $\dot{\delta}$, and $\dot{\mathcal{U}}$ be $S$-names which $d$ forces satisfy the above properties.
Since $S$ is countably distributive, we can extend $d$ to $d^{\prime}$ which decides $\dot{\delta}$ and $\dot{U}\cap T_{\dot{\delta}}^{l}$ as $\delta$ and $\mathcal{Y}_{0}$, and decides sequences $\langle\delta_{n}:n<\omega\rangle$ and $\langle\vec{b}_{n}:n<\omega\rangle$ which witness that $\dot{\mathcal{U}}$ is suitable for $F$. Observe that by absoluteness these sequences satisfy the bullet points of Definition 5.5 (Suitable Generalized Promises) in $V$. Working in $V$ and using the countable distributivity of $S$, define $\mathcal{V}$ with domain $[\delta,\omega_{1})$ so that for all $\delta\leq\zeta<\omega_{1}$, $\mathcal{V}(\zeta)$ is the set of all $\mathcal{Y}\subseteq T_{\zeta}^{l}$ such that for some $d^{*}\geq_{S}d^{\prime}$, $d^{*}\Vdash_{S}\mathcal{Y}=\dot{U}\cap T_{\zeta}^{l}$. Note that $\mathcal{V}(\delta)=\{\mathcal{Y}_{0}\}$, and each $\mathcal{V}(\zeta)$ is countable as $S$ is c.c.c. It is now routine to check that $\mathcal{V}$ is a generalized promise in $V$ which is suitable for $F$.
@@
Let $S$ be a normal Suslin tree which forces that $T$ is Aronszajn. To show that $\mathbb{P}$ forces that $S$ is Suslin, it suffices to show that whenever $(F,\Gamma)$ forces that $\dot{E}$ is a dense open subset of $S$, then there exist $(G,\Sigma)\leq(F,\Gamma)$ and $\delta<\omega_{1}$ such that $(G,\Sigma)\Vdash\dot{E}\subseteq S_{\delta}$.
@@ 9. A Lindelöf Tree With Non-Lindelöf Square
For showing that the subspace topology is finer than the fine wedge topology, consider a basic open set $U$ in the fine wedge topology on $S\otimes S$ and we show that $U$ is open in the subspace topology. So $U=(x,y)\!\!\uparrow\setminus F\!\!\uparrow$, for some $(x,y)\in S\otimes S$ and some finite set $F$ of immediate successors of $(x,y)$. To show that $U$ is open in the subspace topology, we consider any $(a,b)\in U$ and find an open set $V$ in the product topology such that $(a,b)\in V\cap(S\otimes S)\subseteq U$. Since $(a,b)\in U$, $x\leq_{S}a$, $y\leq_{S}b$, and for all $(c,d)\in F$, $(a,b)\not\geq_{S\otimes S}(c,d)$.
First, assume that $(a,b)=(x,y)$. Let $X=\{c:\exists d\ (c,d)\in F\}$, which is finite. Define $V=(x\!\!\uparrow\setminus X\!\!\uparrow)\times y\!\!\uparrow$, which is open in the subspace topology on $S\times S$. Clearly, $(x,y)\in V$. We claim that $V\cap(S\otimes S)\subseteq U$. Consider $(e,f)\in V\cap(S\otimes S)$. If $(e,f)=(x,y)$, then $(e,f)\in U$. Otherwise, the height of $e$ and $f$ is greater than the height of $x$ and $y$. So $x<_{S}e$ and $y<_{S}f$, and hence $(e,f)\in(x,y)\!\!\uparrow$. Suppose for a contradiction that $(e,f)\in F\!\!\uparrow$. Fix $(c,d)\in F$ such that $(c,d)\leq_{S\otimes S}(e,f)$. Then $c\in X$ and hence $e\in X\!\!\uparrow$, which contradicts that $(e,f)\in V$.
Secondly, assume that $(a,b)\neq(x,y)$. Then $x<_{S}a$ and $y<_{S}b$. Define $V=a\!\!\uparrow\times b\!\!\uparrow$, which is open in the product topology on $S\times S$. Clearly, $(a,b)\in V$. We claim that $V\cap(S\otimes S)\subseteq U$. Consider $(e,f)\in V\cap(S\otimes S)$. Then $a\leq_{S}e$ and $b\leq_{S}f$, and hence $(e,f)\in(x,y)\!\!\uparrow$. Suppose for a contradiction that $(e,f)\in F\!\!\uparrow$. Fix $(c,d)\in F$ such that $(c,d)\leq_{S\otimes S}(e,f)$, that is, $c\leq_{S}e$ and $d\leq_{S}f$. As the height of $a$ and $b$ is at least the height of $c$ and $d$, it follows that $c\leq_{S}a$ and $d\leq_{S}b$. Hence, $(a,b)\in F\!\!\uparrow$, which contradicts that $(a,b)\in U$.
@@
Obviously, $H$ is surjective. For continuity, consider a basic open set $U$ in the fine wedge topology on $S\otimes S$. Then $U=(x,y)\!\!\uparrow\setminus F\!\!\uparrow$ for some $(x,y)\in S\otimes S$ and some finite set $F$ of immediate successors of $(x,y)$. We claim that $H^{-1}(U)$ is open in $S\times S$. Let $\beta$ be the height of $x$ and $y$.
So let $(a,b)\in H^{-1}(U)$, and we find an open set $V\subseteq S\times S$ such that $(a,b)\in V\subseteq H^{-1}(U)$. Let $\gamma=\min(\mathrm{ht}_{S}(a),\mathrm{ht}_{S}(b))$.
Case 1: $\gamma>\beta$. Then $x<_{S}a\upharpoonright\gamma$ and $y<_{S}b\upharpoonright\gamma$. We claim that $V=a\!\!\uparrow\times\ b\!\!\uparrow$ works. Clearly, $(a,b)\in V$. Suppose $(c,d)\in V$, and we show that $H(c,d)\in U$. Let $\zeta=\min(\mathrm{ht}_{S}(c),\mathrm{ht}_{S}(d))$, so $H(c,d)=(c\upharpoonright\zeta,d\upharpoonright\zeta)$. Then $\beta<\gamma\leq\zeta$. So $x<_{S}a\upharpoonright\gamma\leq_{S}c\upharpoonright\zeta$ and $y<_{S}b\upharpoonright\gamma\leq_{S}d\upharpoonright\zeta$. Therefore, $H(c,d)\in(x,y)\!\!\uparrow$. Suppose for a contradiction that there exists $(w,z)\in F$ such that $(w,z)\leq_{S\otimes S}(c\upharpoonright\zeta,d\upharpoonright\zeta)$. As $(a\upharpoonright\gamma,b\upharpoonright\gamma)\leq_{S\otimes S}(c\upharpoonright\zeta,d\upharpoonright\zeta)$, it follows that $(a\upharpoonright\gamma,b\upharpoonright\gamma)$ and $(w,z)$ are comparable in $S\otimes S$. But $w$ and $z$ have height $\beta+1$ and $a\upharpoonright\gamma$ and $b\upharpoonright\gamma$ have height $\gamma\geq\beta+1$. So $(w,z)\leq_{S\otimes S}(a\upharpoonright\gamma,b\upharpoonright\gamma)$ and thus $(a\upharpoonright\gamma,b\upharpoonright\gamma)=H(a,b)\in F\!\!\uparrow$, which is a contradiction.
Case 2: $\gamma=\beta$. So either $a=x$ or $b=y$. By symmetry, without loss of generality assume the former. Let $X=\{w:\exists z\ (w,z)\in F\}$. Then $X$ is finite. We claim that $V=(x\!\!\uparrow\setminus X\!\!\uparrow)\times b\!\!\uparrow$ works. Clearly, $(a,b)\in V$. Let $(e,f)\in V$, and we show that $H(e,f)\in U$. Let $\zeta=\min(\mathrm{ht}_{S}(e),\mathrm{ht}_{S}(f))$. Obviously, $\beta\leq\zeta$. So $x\leq_{S}e\upharpoonright\zeta$ and $y\leq_{S}f\upharpoonright\zeta$. Suppose for a contradiction that $H(e,f)=(e\upharpoonright\zeta,f\upharpoonright\zeta)\in F\!\!\uparrow$. Then $\zeta\geq\beta+1$ and $(e\upharpoonright(\beta+1),f\upharpoonright(\beta+1))\in F$. Then $e\upharpoonright(\beta+1)\in X$ and so $e\in X\!\!\uparrow$. But $(e,f)\in V$ implies that $e\notin X\!\!\uparrow$, which is a contradiction. ∎
@@ 10. Additional Properties of the Forcing Poset
Assuming CH, any forcing poset which has the $\omega_{2}$-p.i.c. is $\omega_{2}$-c.c.
@@
Let $\alpha\leq\omega_{2}$. Suppose that $\langle\mathbb{P}_{i},\dot{\mathbb{Q}}_{j}:i\leq\alpha,\ j<\alpha\rangle$ is a countable support forcing iteration such that for all $i<\alpha$, $\mathbb{P}_{i}$ forces that ˙$\dot{\mathbb{Q}}_{i}$ has the $\omega_{2}$-p.i.c. If $\alpha<\omega_{2}$, then $\mathbb{P}_{\alpha}$ has the $\omega_{2}$-p.i.c., and if $\alpha=\omega_{2}$ and CH holds, then $\mathbb{P}_{\omega_{2}}$ is $\omega_{2}$-c.c.
@@
By Lemma 6.7, $(G,\Sigma\cup h[\Sigma])$ is a condition which extends both $(G,\Sigma)$ and $(G,h[\Sigma])$, and hence is a total master condition for both $N_{\alpha}$ and $N_{\beta}$ which extends $(F,\Gamma)$ and $h(F,\Gamma)$. Now consider $u\in N_{\alpha}\cap\mathbb{P}$ and $r\leq(G,\Sigma\cup h[\Sigma])$. We find $s\leq r$ such that $s\leq u$ iff $s\leq h(u)$. Consider the dense open set $D$ in $N_{\alpha}$ of conditions which are either below $u$ or incompatible with $u$. Fix $v\in D\cap N_{\alpha}$ such that $(G,\Sigma)\leq v$. By the previous paragraph, $(G,h[\Sigma])\leq h(v)$. As $r\leq(G,\Sigma\cup h[\Sigma])\leq(G,\Sigma),(G,h[\Sigma])$, it follows that $r\leq v,h(v)$. Case 1: $v\leq u$. Then $h(v)\leq h(u)$. So $r\leq u,h(u)$. Case 2: $v$ is incompatible with $u$. Then $h(v)$ is incompatible with $h(u)$. Since $r\leq v,h(v)$, $r$ is incompatible with $u$ and $h(u)$, and in particular, $r$ does not extend $u$ or $h(u)$. ∎
@@
Suppose that $\langle\mathbb{P}_{i},\dot{\mathbb{Q}}_{j}:i\leq\alpha,\ j<\alpha\rangle$ is a countable support forcing iteration such that for all $i<\alpha$, $\mathbb{P}_{i}$ forces that ˙$\dot{\mathbb{Q}}_{i}$ is $(<\!\omega_{1})$-proper and Dee-complete. Then $\mathbb{P}_{\alpha}$ does not add reals.
@@ 11. Consistency Results
By Lemma 1.1(a), the assumption implies that every normal infinitely splitting Aronszajn tree contains an uncountable downwards closed $n$-splitting normal subtree. Let $T$ be an Aronszajn tree. We may assume that $T$ has a root, for otherwise we can restrict our attention to $T_{x}$ for some $x$ of height 00. If $T$ is not Hausdorff, then by a standard shifting argument find a Hausdorff tree $T^{+}$ such that $T^{+}$ is equal to $T$ restricted to non-limit levels. Note that if $T^{+}$ has an uncountable downwards closed $(\leq\!m)$-splitting subtree, then so does $T$. So without loss of generality assume that $T=T^{+}$. By Lemma 1.1(a), $T$ contains an uncountable downwards closed subtree $U$ such that every element of $U$ has uncountably many elements of $U$ above it. So again let us just assume that $T=U$. Then $T$ is normal. So by Lemma 1.1(b), there exists a club $C\subseteq\omega_{1}$ such that $T\upharpoonright C$ is infinitely splitting. Note that $T\upharpoonright C$ is normal. Find an uncountable downwards closed normal subtree $W\subseteq T\upharpoonright C$ which is $m$-splitting. By the normality of $W$, it is easy to check that the downwards closure of $W$ in $T$ is $(\leq\!m)$-splitting, and we are done. ∎
@@
Start with a model satisfying GCH. Define a countable support forcing iteration $\langle\mathbb{P}_{i},\dot{\mathbb{Q}}_{j}:i\leq\omega_{2},\ j<\omega_{2}\rangle$ where we bookkeep so that for every $2\leq n<\omega$ and every normal $(\geq\!n)$-splitting normal Aronszajn tree $T$ in the final model, there exists some $i<\omega_{2}$ such that $\Vdash_{i}\dot{\mathbb{Q}}_{i}=\mathbb{P}(T,\{n\})$. By Proposition 10.4 and Theorems 10.2 and 10.3, this bookkeeping is possible since each $\mathbb{P}_{i}$ has size $\omega_{2}$ and is $\omega_{2}$-c.c. By Propositions 10.6 and 10.8 and Theorem 10.7, $\mathbb{P}_{\omega_{2}}$ does not add reals, and in particular, preserves $\omega_{1}$ and CH. We claim that in $V^{\mathbb{P}_{\omega_{2}}}$ there do not exist any Lindelöf trees. Start with an $\omega_{1}$-tree $T$. If $T$ is not Aronszajn, then it is not Lindelöf by [Mar, Lemma 4.2]. So assume that $T$ is Aronszajn. By Lemma 11.1, $T$ contains an uncountable downwards closed $(\leq\!2)$-splitting subtree. By Theorem 2.7, $T$ is not Lindelöf. ∎
@@
Let $2\leq m<\omega$. Suppose that every normal infinitely splitting Aronszajn which does not contain a Suslin subtree contains an uncountable downwards closed $m$-splitting subtree. Then every Aronszajn tree which does not contain a Suslin subtree contains an uncountable downwards closed $(\leq\!m)$-splitting subtree.
@@
Let $2<n\leq\omega$. It is consistent that CH holds, every normal $(\geq\!n)$-splitting Aronszajn tree contains an uncountable downwards closed $n$-splitting normal subtree, and there exists a normal Aronszajn tree which is infinitely splitting but does not contain an uncountable downwards closed $(<\!n)$-splitting subtree.
@@
Start with a model satisfying GCH and there exists a normal infinitely splitting Suslin tree $S$. By [Mar, Theorem 4.12], $S$ does not contain an uncountable downwards closed finitely splitting subtree. In particular, $S$ does not contain an uncountable downwards closed $(<\!n)$-splitting subtree. Now define a countable support forcing iteration $\langle\mathbb{P}_{i},\dot{\mathbb{Q}}_{j}:i\leq\omega_{2},\ j<\omega_{2}\rangle$ as in the proof of Theorem 11.2, except that we only force with $\mathbb{P}(T,\{n\})$ for normal Aronszajn trees $T$ which are $(\geq\!n)$-splitting. By Theorem 7.1, each such forcing preserves the fact that $S$ does not contain an uncountable downwards closed $(<\!n)$-splitting subtree, so by Theorem 3.1 so does $\mathbb{P}_{\omega_{2}}$. ∎

### v2

@@ 7. Not Adding Small-Splitting Subtrees
Let $1<n\leq s$ and suppose that $S$ is an $\omega_{1}$-tree which does not contain an uncountable downwards closed (<n)(<\!n)-splitting subtree. Consider a condition $(F,\Gamma)\in\mathbb{P}$ with some top level $\beta$ and assume that $(F,\Gamma)$ forces that $\dot{W}$ is a downwards closed (<n)(<\!n)-splitting subtree of $S$. We find an extension of $(F,\Gamma)$ which forces that $\dot{W}$ is countable.
@@ 8. Preserving Suslin Trees
Define $B$ as the set of all $b\in S$ satisfying that $c\upharpoonright\beta<_{S}b$, and letting $\delta=\mathrm{ht}_{S}(b)$, there exists some $\vec{y}=(y_{0},\ldots,y_{l-1})\in T_{\vec{x}}\cap T_{\delta}^{l}$ such that whenever $(G,\Sigma)\leq(F,\Gamma)$ has some top level $\gamma\leq\delta$ and for all $i<l$, $G(y_{i}\upharpoonright\gamma)=1$, then it is not the case that $(G,\Sigma)\Vdash b\in\dot{E}$.
@@
These properties imply that $B$ is uncountable and downwards closed in $S_{c\upharpoonright\beta}$.
For any $e\in B$, let $\delta_{e}=\mathrm{ht}_{S}(e)$ and define $\mathcal{Y}_{e}$ to be the set of all $\vec{y}=(y_{0},\ldots,y_{l-1})\in T_{\vec{x}}\cap T_{\delta_{e}}^{l}$ witnessing that $e\in B$, which means that whenever $(G,\Sigma)\leq(F,\Gamma)$ has some top level $\gamma\leq\delta_{e}$ and for all $i<l$, $G(y_{i}\upharpoonright\gamma)=1$, then it is not the case that $(G,\Sigma)\Vdash e\in\dot{E}$. Note that by (3) above, for all $e$ and $f$ in $B$ such that $e<_{S}f$, $\mathcal{Y}_{f}\upharpoonright\delta_{e}\subseteq\mathcal{Y}_{e}$.
@@
By our assumption on $S$, $T$ is still Aronszajn in $V[G_{S}]$. So we can apply Lemma 5.6 in $V[G_{S}]$ to find a countable limit ordinal $\delta>\beta$ such that the function $\mathcal{U}$ with domain $\delta,\omega_{1})$ defined by $\mathcal{U}(\gamma)=\{U\cap T_{\gamma}^{l}\}$ is a generalized promise which is suitable for $F$. Now let $\dot{W}$, $\dot{U}$, $\dot{\delta}$, and $\dot{\mathcal{U}}$ be $S$-names which $d$ forces satisfy the above properties.
Since $S$ is countably distributive, we can extend $d$ to $d^{\prime}$ which decides $\dot{\delta}$ and $\dot{U}\cap T_{\dot{\delta}}^{l}$ as $\delta$ and $\mathcal{Y}_{0}$, and decides sequences $\langle\delta_{n}:n<\omega\rangle$ and $\langle\vec{b}_{n}:n<\omega\rangle$ which witness that $\dot{\mathcal{U}}$ is suitable for $F$. Observe that by absoluteness these sequences satisfy the bullet points of Definition 5.5 (Suitable Generalized Promises) in $V$. Working in $V$ and using the countable distributivity of $S$, define $\mathcal{V}$ with domain $\delta,\omega_{1})$ so that for all $\delta\leq\zeta<\omega_{1}$, $\mathcal{V}(\zeta)$ is the set of all $\mathcal{Y}\subseteq T_{\zeta}^{l}$ such that for some $d^{*}\geq_{S}d^{\prime}$, $d^{*}\Vdash_{S}\mathcal{Y}=\dot{U}\cap T_{\zeta}^{l}$. Note that $\mathcal{V}(\delta)=\{\mathcal{Y}_{0}\}$, and each $\mathcal{V}(\zeta)$ is countable as $S$ is c.c.c. It is now routine to check that $\mathcal{V}$ is a generalized promise in $V$ which is suitable for $F$.
@@
Let $S$ be a normal Suslin tree which forces that $T$ is Aronszajn. To show that $\mathbb{P}$ forces that $S$ is Suslin, by Lemma 1.1(d) it suffices to show that whenever $(F,\Gamma)$ forces that $\dot{E}$ is a dense open subset of $S$, then there exist $(G,\Sigma)\leq(F,\Gamma)$ and $\delta<\omega_{1}$ such that $(G,\Sigma)\Vdash S_{\delta}\subseteq\dot{E}$.
@@ 9. A Lindelöf Tree With Non-Lindelöf Square
For showing that the subspace topology is finer than the fine wedge topology, it suffices to show that every member of the subbase of the fine wedge topology on $S\otimes S$ is open in the subspace topology. Consider $U=(x,y){\uparrow}$ for some $(x,y)\in S\otimes S$. Then $U=((x{\uparrow})\times(y{\uparrow}))\cap(S\otimes S)$, which is open in the subspace topology. Now consider $V=S\otimes S\setminus(x,y){\uparrow}$ for some $(x,y)\in S\otimes S$. Let $(a,b)\in V$, and we will find a set $W$ which is open in $S\times S$ such that $(a,b)\in W\cap(S\otimes S)\subseteq V$. By the definition of $V$, either $a\not\geq_{T}x$ or $b\not\geq_{T}y$. By symmetry, assume the former. Now check that $W=(T\setminus x{\uparrow})\times b{\uparrow}$ works.
@@
Clearly, $H$ is surjective. For continuity, it suffices to show that whenever $U$ is a member of the subbase of the fine wedge topology on $S\otimes S$, then $H^{-1}(U)$ is open in $S\times S$. First, suppose that $U=(x,y){\uparrow}$, where $(x,y)\in S\otimes S$. It is straightforward to show that $H^{-1}(U)=x{\uparrow}\times y{\uparrow}$, which is open in $S\times S$. Secondly, assume that $V=S\otimes S\setminus(x,y){\uparrow}$, where $(x,y)\in S\otimes S$. Then $H^{-1}(V)=S\times S\setminus H^{-1}((x,y){\uparrow})=S\times S\setminus(x{\uparrow}\times y{\uparrow})$, which is also open in $S\times S$. ∎
@@ 10. Additional Properties of the Forcing Poset
Assuming CH, any forcing poset which has the $\omega_{2}$-p.i.c. is $\omega_{2}$-$c.$
@@
Let $\alpha\leq\omega_{2}$. Suppose that $\langle\mathbb{P}_{i},\dot{\mathbb{Q}}_{j}:i\leq\alpha,\ j<\alpha\rangle$ is a countable support forcing iteration such that for all $\alpha, i\mathbb{P}_{i}$ forces that $\dot{\mathbb{Q}}_{i}$ has the $\omega_{2}$-p.i.c. If $\alpha<\omega_{2}$, then $\mathbb{P}_{\alpha}$ has the $\omega_{2}$-p.i.c., and if $\alpha=\omega_{2}$ and CH holds, then $\mathbb{P}_{\omega_{2}}$ is $\omega_{2}$-$c.$
@@
By Lemma 6.7, $(G,\Sigma\cup h[\Sigma])$ is a condition which extends both $(G,\Sigma)$ and $(G,h[\Sigma])$, and hence is a total master condition for both $N_{\alpha}$ and $N_{\beta}$ which extends $(F,\Gamma)$ and $h(F,\Gamma)$. Now consider $u\in N_{\alpha}\cap\mathbb{P}$ and $r\leq(G,\Sigma\cup h[\Sigma])$. We find $s\leq r$ such that $s\leq u$ iff $s\leq h(u)$. Consider the dense open set $D$ in $N_{\alpha}$ of conditions which are either below $u$ or incompatible with $u$. Fix $v\in D\cap N_{\alpha}$ such that $(G,\Sigma)\leq v$. By the previous paragraph, $(G,h[\Sigma])\leq h(v)$. As $r\leq(G,\Sigma\cup h[\Sigma])\leq(G,\Sigma),(G,h[\Sigma])$, it follows that $r\leq v,h(v)$. Case 1: $v\leq u$. Then $h(v)\leq h(u)$. So $r\leq u,h(u)$. Case 2: $v$ is incompatible with $u$. Then $h(v)$ is incompatible with $h(u)$. Since $\leq v,h(v), rr$ is incompatible with $u$ and $h(u)$, and in particular, $r$ does not extend $u$ or $h(u)$. ∎
@@
Suppose that $\langle\mathbb{P}_{i},\dot{\mathbb{Q}}_{j}:i\leq\alpha,\ j<\alpha\rangle$ is a countable support forcing iteration such that for all $\alpha, i\mathbb{P}_{i}$ forces that $\dot{\mathbb{Q}}_{i}$ is $(<\!\omega_{1})$-proper and Dee-complete. Then $\mathbb{P}_{\alpha}$ does not add reals.
@@ 11. Consistency Results
By Lemma 1.1(a), the assumption implies that every normal infinitely splitting Aronszajn tree contains an uncountable downwards closed $n$-splitting normal subtree. Let $T$ be an Aronszajn tree. We may assume that $T$ has a root, for otherwise we can restrict our attention to $T_{x}$ for some $x$ of height 0. If $T$ is not Hausdorff, then by a standard shifting argument find a Hausdorff tree $T^{+}$ such that $T^{+}$ is equal to $T$ restricted to non-limit levels. Note that if $T^{+}$ has an uncountable downwards closed $(\leq\!m)$-splitting subtree, then so does $T$. So without loss of generality assume that $T=T^{+}$. By Lemma 1.1(a), $T$ contains an uncountable downwards closed subtree $U$ such that every element of $U$ has uncountably many elements of $U$ above it. So again let us just assume that $T=U$. Then $T$ is normal. So by Lemma 1.1(b), there exists a club $C\subseteq\omega_{1}$ such that $T\upharpoonright C$ is infinitely splitting. Note that $T\upharpoonright C$ is normal. Find an uncountable downwards closed normal subtree $W\subseteq T\upharpoonright C$ which is $m$-splitting. By the normality of $W$, it is easy to check that the downwards closure of $W$ in $T$ is $(\leq\!m)$-splitting, and we are done. ∎
@@
Start with a model satisfying GCH. Define a countable support forcing iteration $\langle\mathbb{P}_{i},\dot{\mathbb{Q}}_{j}:i\leq\omega_{2},\ j<\omega_{2}\rangle$ where we bookkeep so that for every $2\leq n<\omega$ and every normal $(\geq\!n)$-splitting normal Aronszajn tree $T$ in the final model, there exists some $i<\omega_{2}$ such that $\Vdash_{i}\dot{\mathbb{Q}}_{i}=\mathbb{P}(T,\{n\})$. By Proposition 10.4 and Theorems 10.2 and 10.3, this bookkeeping is possible since each $\mathbb{P}_{i}$ has size $\omega_{2}$ and is $\omega_{2}$-c.c. By Propositions 10.6 and 10.8 and Theorem 10.7, $\mathbb{P}_{\omega_{2}}$ does not add reals, and in particular, preserves $\omega_{1}$ and CH. We claim that in $V^{\mathbb{P}_{\omega_{2}}}$ there do not exist any Lindelöf trees. Start with an $\omega_{1}$-tree $T$. If $T$ is not Aronszajn, then it is not Lindelöf by [Mar24b, Lemma 4.2]. So assume that $T$ is Aronszajn. By Lemma 11.1, $T$ contains an uncountable downwards closed $(\leq\!2)$-splitting subtree. By Theorem 2.7, $T$ is not Lindelöf. ∎
@@
Let $2\leq m<\omega$. Suppose that every normal infinitely splitting Aronszajn tree which does not contain a Suslin subtree contains an uncountable downwards closed $m$-splitting subtree. Then every Aronszajn tree which does not contain a Suslin subtree contains an uncountable downwards closed $(\leq\!m)$-splitting subtree.
@@
Let $2<n\leq\omega$. It is consistent that CH holds, every normal $(\geq\!n)$-splitting Aronszajn tree contains an uncountable downwards closed $n$-splitting normal subtree, and there exists a normal Aronszajn tree which is infinitely splitting but does not contain an uncountable downwards closed (<n)(<\!n)-splitting subtree.
@@
Start with a model satisfying GCH and there exists a normal infinitely splitting Suslin tree $S$. By [Mar24b, Theorem 4.12], $S$ does not contain an uncountable downwards closed finitely splitting subtree. In particular, $S$ does not contain an uncountable downwards closed (<n)(<\!n)-splitting subtree. Now define a countable support forcing iteration $\langle\mathbb{P}_{i},\dot{\mathbb{Q}}_{j}:i\leq\omega_{2},\ j<\omega_{2}\rangle$ as in the proof of Theorem 11.2, except that we only force with $\mathbb{P}(T,\{n\})$ for normal Aronszajn trees $T$ which are $(\geq\!n)$-splitting. By Theorem 7.1, each such forcing preserves the fact that $S$ does not contain an uncountable downwards closed (<n)(<\!n)-splitting subtree, so by Theorem 3.1 so does $\mathbb{P}_{\omega_{2}}$. ∎
Acknowledgements: The author thanks Pedro Marun for providing helpful comments and corrections on an earlier draft of the article, and the referee for writing a careful report.
Funding: The author acknowledges support from the Simons Foundation under the Travel Support for Mathematicians gift 631279.